OpenStudy (anonymous):

write the equation of the line in slope-intercept form, and graph the line. slope= 2/3, y-intercept= -4

OpenStudy (anonymous):

y=mx+b y=2/3+4 you"ll have to graph it yourself its really not hard promise

OpenStudy (anonymous):

made a typo y=2/3x+4

OpenStudy (anonymous):

and the 4 goes on the y axis then you just go up 2 and over 3?

OpenStudy (anonymous):

yess the -4 hunn

OpenStudy (anonymous):

thank you! (:

OpenStudy (anonymous):

No problem :))